Fermi National Laboratory


Accelerator Update for April 27, 2005 - April 29, 2005

Wednesday April 27
The day shift began with Operations monitoring the Recycler stash, with Tevatron (TeV) experts conducting studies, with the TeV separators in a conditioning mode, with cryo system technicians working on wet engines, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and Meson MCenter experiments taking beam.

Pbar stopped stacking at 8:38 AM.

Operations made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 9 AM.

Cryo techs completed their engine work at 11:09 AM.

CDF came out of access at 11:45 AM.

Separator conditioning ended at 1:18 PM.

D0 came out of access at 1:31 PM.

Operations put the TeV through a dry squeeze at 2:54 PM.

Operations put the TeV into shot setup at 3:21 PM

At 4:51 PM, the Booster suffered a RF watchdog trip. Operators reset it and then it tripped again. Operators turned off MiniBooNE beam while loading the TeV with protons.

Booster experts identified a quadrupole damper at 5:02 PM as the cause of the RF trips. They made adjustments and then Operations turned MiniBooNE back on.

At 5:29 PM, Operations reported that they were still having Booster RF problems.

Operations established combination store 4116 at 6:51 PM with an initial luminosity of 119.03E30. This store set a NEW LUMINOSITY RECORD. Pbar resumed stacking.

NuMI experts began target scans at 8:53 PM. A Recycler expert also began a study period.

At 9:18 PM, a Booster power supply (EAPS) tripped off. Operators reset the power supply, but it tripped off again. An expert suggested they turn off BRF19 until NuMI stops running. This resolved the problem.

NuMI target scans ended at 11:13 PM.

At 11:23 PM, CDF reported a strange spike in the beam. Operators investigated and found that TRF2 had sparked, but didn't trip off. D0 saw a small drop in its luminosity.

At 11:55 PM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum problem. They both reset okay.

Thursday April 28
The midnight shift began with Operations monitoring store 4116, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the Recycler into shot setup to stash antiprotons at 1:02 AM.

At 2:06 AM, TRF2 sparked again.

The Recycler stash ended at 2:22 AM. Pbar resumed stacking.

MCenter resumed taking beam at 2:36 AM.

A Pbar expert worked on improving the stacking rate until 5 AM. Both antiproton production and stacking were improved.

The Recycler began stashing antiprotons at 9:05 AM.

The stashing ended and Pbar resumed stacking at 10:20 AM.

A Pbar expert began a reverse proton study at 12:10 PM.

At 1:01 PM, BRF16 tripped multiple times. It reset okay but then tripped again three minutes later. It reset okay.

At 1:35 PM, a Linac RF station (KRF5) tripped off. Operators investigated and managed to get the station back up by 1:43 PM.

NuMI experts began a target scan at 6:29 PM.

Friday April 29
The midnight shift began with Operations monitoring store 4116 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:44 AM, Operations reported an alarm on the TeV sector C4 QPM. An operator disabled the alarm and will reboot the unit after the store gets terminated.

Operations terminated store 4116 at 3:04 AM.

At 3:46 AM, Operations reported another problem with the C4 QPM. It wouldn't reboot. An operator reset the QPM locally, which caused the TeV ramp to dump, but the QPM did reset.

At 4:11 AM, Operations reported that the C4 QPM showed a quench, but the TeV hadn't quenched. Operators contacted an expert who came in to check the hardware.

Since the TeV was being held off due to the QPM problem, operators put Pbar back into a stacking mode at 5:42 AM.

The QPM expert reported at 5:48 AM that he was replacing the control card. At 6:16 AM, the system was still showing a quench.

Operations disabled beam to MiniBooNE at 5:57 AM for scheduled work.

MCenter resumed taking beam at 6:25 AM.

The QPM expert reported at 6:33 AM that the quench indication at C4 was true. Operations, with expert help, managed to reset the status. Unfortunately, the C4 power supply would not reset due to a quench and tripped breaker indication.

At 6:49 AM, operators investigated the C4 power supply; they found no problems. The expert investigated further and at 7:10 AM discovered and fixed a bad connection between the C4 QPM and the power supply.

Operations put the TeV into a dry squeeze at 7:25 AM.

Operations put the TeV into shot setup at 7:49 AM.

The Plan for Friday and the Weekend
The plan for today is to establish a store this morning and then for the rest of the weekend to stack, stash and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/29/2005   email Fermilab